### v2

# Position: From Correlation to Causation: Max-Pooling-Based Multi-Instance Learning Leads to More Robust Whole Slide Image Classification

Abstract
Although attention-based multi-instance learning (MIL) algorithms have achieved impressive performance on slide-level whole slide image (WSI) classification tasks, they are prone to mistakenly focusing on irrelevant patterns such as staining conditions and tissue morphology, leading to incorrect patch-level predictions and unreliable interpretability. In this paper, we analyze why attention-based methods tend to rely on spurious correlations in their predictions. Furthermore, we revisit max-pooling-based approaches and examine the reasons behind the underperformance of existing methods. We argue that well-trained max-pooling-based MIL models can make predictions based on causal factors and avoid relying on spurious correlations. Building on these insights, we propose a simple yet effective max-pooling-based MIL method (FocusMIL) that outperforms existing mainstream attention-based methods on two datasets. In this position paper, we advocate renewed attention to max-pooling-based methods to achieve more robust and interpretable predictions.
Multi-Instance Learning, Reproducibility
1 Introduction
Figure 1: (a) A causal graph illustrating the generative procedure of patches and their labels. Ideally, $\bm{Z}_{C}$ is used to predict $y$. (b) The architecture of attention-based MIL (excluding the bag feature for brevity), where the classifier can leverage any of $\bm{Z}_{C}$, $\bm{Z}_{S}$, and/or $\bm{Z}_{B}$ for prediction.
Whole slide images (WSIs) analysis has been widely utilized in clinical applications such as computer-aided tumour diagnosis and prognosis (Lu et al., 2021; Yao et al., 2020; Chen et al., 2021; Li et al., 2021). As a single WSI often contains billions of pixels and obtaining patch-level annotations is labour-intensive (Lu et al., 2021), it is a standard practice to crop the WSI into smaller patches and utilize Multi-Instance Learning (MIL) for classification. Multi-Instance Learning is a weakly-supervised learning paradigm designed to learn from samples represented as bags (slides) of instances (patches). As computer-aided diagnosis models must offer interpretable predictions at both the slide and patch levels to support clinical decision-making, MIL is particularly well-suited (Li et al., 2021; Qu et al., 2022).
Most existing MIL-based WSI classification algorithms focus on slide-level predictions, using attention (Vaswani et al., 2017) for aggregating patch-level representations (Ilse et al., 2018; Li et al., 2021; Shao et al., 2021; Qu et al., 2022; Zhang et al., 2022a, 2023; Tang et al., 2023, 2024; Keum et al., 2024). As Figure 1 shows, they assign attention weights to instances within a bag and aggregate their weighted instance representations to a bag representation for classification. However, recent studies (Raff & Holt, 2023) have shown that most attention-based MIL methods deviate from the standard MIL assumption, and are not learnable at the instance-level (Jang & Kwon, ).
The standard MIL assumption states that a multi-instance bag is positive if and only if at least one of its instances is positive (Foulds & Frank, 2010; Dietterich et al., 1997). This assumption aligns well with the nature of cancer in WSI, where a slide is considered malignant if cancerous tissue or cells are present, even in a small region. Unfortunately, attention-based bag-level MIL methods can gather information from all the instances in the bag to make holistic discrimination, thereby diverging from this assumption. The violations of this assumption by attention-based methods may limit their interpretability and trustworthiness in medical diagnosis (Raff & Holt, 2023).
Figure 1(a) illustrates a plausible causal structure underlying the generation process of a patch $\bm{x}$, its ground-truth label $y$, and its predicted label $\hat{y}$. We posit that patch features in WSIs are determined by three sets of factors: $\bm{Z}_{c}$, which causally influence whether a cell is cancerous, such as cellular morphology; $\bm{Z}_{s}$, which dictate stylistic elements, such as the positioning, size, and compression state of cells within the patch; and $\bm{Z}_{B}$, which are derived from the whole slide, including tissue structures, hematoxylin and eosin staining conditions, and additional biases or artifacts introduced during slide preparation. Although an ideal model should make predictions based on $\bm{Z}_{c}$, most current models do not distinguish between $\bm{Z}_{c}$ and $\bm{Z}_{B}$.
Attention-based models form bag-level features by weighting instances within each bag, as shown in Figure 1(b). This process can inadvertently leverage the bias-related factor $\bm{Z}_{B}$ to fit the training set by aggregating the instances within a bag with weighted attention rather than focusing on the true causal indicator $\bm{Z}_{C}$. As a result, these models can misinterpret non-causal or irrelevant instances as influential and produce incorrect predictions when the biases in unseen slides differ from those in the training set, undermining both robustness and interoperability. Moreover, attention-based models often miss tumour cells that are morphologically similar to normal cells and less malignant (Jögi et al., 2012; Zhang et al., 2023; Tang et al., 2023), leading to potential missed diagnoses in early-stage cancer. Their tendency to prioritize salient instances (Tang et al., 2023; Zhang et al., 2023) and rely on spurious correlations to fit the more challenging slides during training exacerbates these issues.
Although max-pooling-based methods inherently adhere to the standard MIL assumption (Raff & Holt, 2023) and have been theoretically shown to be learnable at the instance level (Jang & Kwon, ), they have been largely overlooked in recent years due to their relatively weaker classification performance. This naturally raises the question: why do max-pooling methods, which appear well-suited for WSI classification (where the presence of a single cancerous instance determines the bag label), perform poorly? Moreover, can they be refined to leverage their inherent advantages for more robust and interpretable WSI classification?
This position paper starts by empirically demonstrating that max-pooling-based methods can sometimes fail to learn discriminative features when the positive concept is difficult to capture, often overfitting the training set by memorizing the instances within the bag and performing poorly on the test set. This issue is particularly pronounced when max-pooling is combined with a feed-forward neural network. Motivated by these insights, we propose FocusMIL, a simple yet effective max-pooling-based approach that encodes latent variables into a multivariate Gaussian distribution to mitigate naive memorization. FocusMIL integrates a feed-forward neural network, reparameterization operations, and a max-pooling operator. The key differences of our method from traditional max-pooling-based MIL classification methods are: 1) learning a distribution over latent representations rather than a deterministic one, and 2) using mini-batch gradient descent with a batch size greater than one instead of a batch size of one. We discuss how the latent distribution and a larger batch enable the model to better recognize hard instances and disentangle $\bm{Z}_{c}$ and $\bm{Z}_{B}$.
We validate our position with comprehensive evaluations of attention-based and max-pooling-based algorithms on the Camelyon16 (Bejnordi et al., 2017a) and Camelyon17 (Bejnordi et al., 2017b) datasets. The results show that the max-pooling-based FocusMIL achieves comparable or better performance than state-of-the-art attention-based MIL methods at the slide-level, while significantly outperforming attention-based methods in patch-level predictions and tumour region localization. Furthermore, our experimental results on the semi-synthetic dataset confirm that the performance of attention-based methods degrades significantly when the bag context distribution differs substantially between the training and test sets, whereas max-pooling-based methods can still achieve strong results. Therefore, we argue for a renewed focus on max-pooling-based methods, emphasizing that they can offer not only improved predictive performance but also more robust and interpretable outputs for real-world clinical applications. The key insights of this paper are summarized as follows:
  1. 1.
    We discuss why attention-based methods are prone to relying on spurious correlations, undermining the reliability of their prediction.
  2. 2.
    We investigate traditional max-pooling-based methods and empirically analyze why they often fail to deliver competitive performance in WSI classification.
  3. 3.
    We propose FocusMIL, a simple max-pooling-based method that outperforms complex attention-based MIL methods, highlighting the need to revisit max-pooling-based approaches.
  4. 4.
    We conduct comprehensive experiments to demonstrate that max-pooling-based methods can learn causal factors, resulting in improved robustness for WSI classification tasks. The code will be made publicly available after acceptance.
2 Related Work
2.1 Attention-based MIL for WSI classification
Most MIL-based WSI classification algorithms aggregate patch-level features into bag-level features according to their attention scores. An early bag-based method, ABMIL (Ilse et al., 2018), utilizes a trainable network to calculate the attention scores of instance features for bag-level classification and obtains a bag-level representation through their weighted sum. Subsequent methods have inherited similar ideas and extensively incorporated various attention mechanisms (Li et al., 2021; Lu et al., 2021; Shao et al., 2021; Zhang et al., 2022a).
Attention-based MIL methods face the challenge of overfitting from three sources (Zhang et al., 2022a; Lin et al., 2023; Tang et al., 2023; Zhang et al., 2023). Firstly, many cancerous slides contain salient instances, i.e., areas with highly differentiated cancer cells that significantly differ from the normal ones (Qu et al., 2024; Jögi et al., 2012). Secondly, other slides may have hard instances, e.g., small cancerous areas where the cancer cells closely resemble normal cells (Bejnordi et al., 2017a). Thirdly, the digital scanning equipment, slide preparation, and staining processes often introduce staining bias into WSI datasets (Zhang et al., 2022c). Existing studies (Tang et al., 2023; Zhang et al., 2023; Keum et al., 2024) have indicated that attention-based WSI classification methods may excessively focus on salient instances and fail to identify hard instances. For positive slides containing only hard instances, these methods often rely on spurious correlations caused by staining bias to fit the training set (Lin et al., 2023). Recent methods like WENO (Qu et al., 2022), MHIM (Tang et al., 2023), IBMIL (Lin et al., 2023), ACMIL (Zhang et al., 2023) and Slot-MIL (Keum et al., 2024) try to mitigate these problems; however, their generalization capability and patch-level classification performances are still not ideal.
2.2 Reproducibility and Learnability in Multiple Instance Learning
Recent studies (Raff & Holt, 2023) have shown that many widely used MIL methods, such as MI-Net (Wang et al., 2018) and TransMIL (Shao et al., 2021), do not adhere to the standard MIL assumptions and may leverage inductive biases, such as the absence of certain instances, as signals for predicting positive bags. As staining bias is prevalent in WSI analysis (Lin et al., 2023; Zhang et al., 2022c), models that do not respect the standard MIL assumptions can learn to use these spurious correlations (Raff & Holt, 2023). Kaczmarzyk et al. (2024) found that attention-based methods heavily rely on non-tumour regions for metastasis detection and often neglect small tumour regions.
mi-Net (Zhou & Zhang, 2002; Wang et al., 2018), and the recently proposed CausalMIL (Zhang et al., 2022b) are perhaps the only two methods that respect the MIL assumptions. CausalMIL applies a variational autoencoder (Kingma & Welling, 2013) to the MIL problem, using a non-factorized prior distribution conditioned on bag information to provide a theoretical guarantee for identifying latent causal representation. However, mi-Net does not perform optimally in WSI classification, and CausalMIL has not yet been studied for WSI classification.
Furthermore, the latest research on the learnability (Jang & Kwon, ) of MIL has theoretically analyzed that all attention-based methods are unlearnable at the instance-level, whereas instance-based approaches are learnable. This further highlights the need for cautious use of attention-based MIL methods and advocates for the exploration of max-pooling-based approaches.
3 Preliminaries
3.1 Multi-Instance Learning(MIL)
A WSI dataset with $n$ slides are treated as MIL bags =$\mathcal{B}=\{\mathbf{B}_{1},\cdots,\mathbf{B}_{i},\cdots,\mathbf{B}_{n}\}$, where each slide has a slide-level label $Y_{i}\in\{0,1\}$ during training. Each slide is then cropped into patches corresponding to MIL instances $\mathbf{B}_{i}=\{\bm{p}_{i1},\cdots,\bm{p}_{ij},\cdots,\bm{p}_{in_{i}}\}$, where $n_{i}$ is the number of patches in the slide. For each patch $\bm{p}_{ij}$, there exists a patch-level label $y_{ij}\in\{0,1\}$ that is unknown to the learner. The standard multi-instance assumption states that a bag is positive if and only if at least one of its instances is positive:
$Y_{i}=\begin{cases}0,&\text{if }\sum_{j}y_{i,j}=0,\\ 1,&\text{else},\end{cases}$ (1)
which is equivalent to using max-pooling on the instances within a bag:
$Y_{i}=\max_{j}\{y_{i,j}\}.$ (2)
Given a feature extractor $f$, each instance is projected onto a $d$-dimensional feature vector $\bm{x}_{ij}=f(\bm{p}_{ij})\in\mathbb{R}^{d\times 1}$.
3.2 Attention-Based MIL
We first briefly review attention-based methods in preparation for our subsequent analyses. These methods aim to learn a weighted aggregation of instance features so that the final bag representation highlights the most relevant patches. A pioneering work in this direction is the ABMIL framework (Ilse et al., 2018).
ABMIL computes an attention weight $\alpha_{ij}$ for each instance $\bm{x}_{ij}$ using a small neural network:
$\alpha_{ij}=\frac{\exp\Bigl{(}\mathbf{w}^{\top}\,\tanh\bigl{(}\mathbf{V}\,\bm{x}_{ij}^{\top}\bigr{)}\Bigr{)}}{\sum_{k=1}^{n_{i}}\exp\Bigl{(}\mathbf{w}^{\top}\,\tanh\bigl{(}\mathbf{V}\,\bm{x}_{ik}^{\top}\bigr{)}\Bigr{)}},$ (3)
where $\mathbf{V}$ and $\mathbf{w}$ are trainable parameters.
The bag-level feature $\bm{z}_{i}$ is then obtained by a weighted sum of instance embeddings:
$\bm{z}_{i}=\sum_{j=1}^{n_{i}}\alpha_{ij}\,\bm{x}_{ij}.$ (4)
Finally, $\bm{z}_{i}$ is fed into a classifier to predict the bag label $\hat{Y}_{i}$. This allows the model to automatically attend to the most “informative” instances in each bag.
Because it effectively integrates instance-level representations into bag-level ones, attention mechanisms have been widely adopted in numerous WSI classification algorithms (Li et al., 2021; Lu et al., 2021; Shao et al., 2021; Zhang et al., 2022a; Qu et al., 2022).
3.3 Max-Pooling-Based MIL
Max-pooling-based MIL enforces the standard MIL assumption by selecting the most “positive” instance within a bag. A representative method, mi-Net, uses a feed-forward neural network followed by a max-pooling operation (Wang et al., 2018). Given instance-level prediction scores $s$ for a bag, the bag prediction is the maximum of the instance scores:
$\hat{Y_{i}}=\max_{j}\{s_{ij}\}.$ (5)
This simple pooling operation guarantees that a bag is classified as positive if it contains at least one positive instance and as negative only if all instances are negative, aligning with the fundamental requirement of WSI classification.
4 Deep MIL for WSI Classification
4.1 Challenges
For WSI classification tasks, slide-level and patch-level predictions should ideally be based on the causal factor $\bm{Z}_{c}$. In practice, however, several challenges may arise:
  1. 1.
    Low proportion of positive patches. In WSI datasets (e.g., Camelyon16 (Bejnordi et al., 2017a)), the percentage of positive patches can be below 10%, making it difficult for models to learn the positive concept.
    Figure 2: Comparison of tumour patches from different tumour slides and normal patches. Tumour cells from slides with small cancerous areas closely resemble normal cells. Conversely, tumour cells from slides with large cancerous areas exhibit very distinct tumour characteristics. Images are from Camelyon16 dataset.
  2. 2.
    Variation in tumour cell morphology. For WSI classification, the morphology of tumour cells may be different across slides due to the varying malignancy levels of cancer in patients (Zhang et al., 2023; Qu et al., 2024; Jögi et al., 2012). As shown in Figure 2, some slides contain large tumour areas with significant morphological differences from normal cells. In contrast, other slides have smaller tumour areas where the tumour cells have higher differentiation and less pronounced differences from normal cells (Qu et al., 2024), which are referred to as “hard instances” in some literature (Tang et al., 2023).
  3. 3.
    Biases introduced during slide preparation. As the slice preparation process may be different across hospitals, the bag context information $\bm{Z}_{B}$ can vary significantly and may correlate with slide labels. For real-world deployment, models should exhibit strong out-of-distribution (OOD) generalization performance.
Next, we discuss how these challenges affect model behaviour in the following two subsections.
4.2 Why Attention-Based MIL Methods are Not Stable
Attention-based MIL methods inherently violate the standard MIL assumption, as the aggregated bag representation freely utilizes information from all instances within a bag (Raff & Holt, 2023). When the morphology of the tumour cells in the slide is more significant and the area is large, these models easily learn discriminative features. However, when tumour cells (hard instances) resemble normal cells, they often struggle to learn the positive concept and instead rely on $\bm{Z}_{B}$ to fit the training set, assigning high attention scores to irrelevant patches. Our visualization results in the Appendix demonstrate that these models may assign high attention scores to “blurred regions”.
Another pitfall of attention-based models prediction is that they may treat $\bm{Z}_{B}$ information from normal tissue as a signal indicating negative slide labels, leading to missed diagnoses. In our experiments on a semi-synthetic dataset, we inject poisons into normal slides in the training set, whereas in the validation set only tumour slides would contain these harmful signals. The results show that the attention-based approaches regard these poisons as a signal of negative slides, and fail entirely on the validation set. Kaczmarzyk et al. (2024) also found that compared to predicting using all patches in a slide, predicting using only tumour patches can achieve higher sensitivity, confirming that ABMIL leverages $\bm{Z}_{B}$ information from normal patches as negative signals. According to the standard MIL assumption, negative instances should not influence the bag label (Raff & Holt, 2023).
In addition, violating the standard multi-instance learning assumptions makes slide-level predictions sensitive to the number of tumour patches. Kaczmarzyk et al. (2024) demonstrated that even when using features extracted by foundation models like UNI (Chen et al., 2024), REMEDIS (Azizi et al., 2023), and CTransPath (Wang et al., 2022), the slide-level predictions are still affected by the size of the tumour regions. This could potentially lead to the disastrous consequence of missed diagnoses.
4.3 Revisiting Max-Pooling-Based MIL
Max-pooling-based MIL inherently aligns the standard multiple-example learning assumption (Dietterich et al., 1997; Raff & Holt, 2023). For WSI classification task, the main goal is to train a patch-level classifier from slide-level supervision while maintaining the standard multi-instance learning assumption and focusing on the causal factor $\bm{Z}_{c}$ that is invariant across slides. Formally, we assume
Assumption 4.1.
The mechanism of determining whether a cell is cancerous from the causal factor $\bm{Z}_{c}$ in patches is invariant across slides.
Corresponding to Assumption 4.1, $\bm{Z}_{s}$ may vary among different patches, while $\bm{Z}_{B}$ may vary among slides.
The Importance of the Standard MIL Assumption.
Models that respect the MIL assumptions predict bag labels solely based on the positive concept, e.g., the presence of tumour cells  (Foulds & Frank, 2010; Dietterich et al., 1997). We argue that max-pooling effectively focuses the model on the causal factor $\bm{Z}_{c}$ for predicting the instance label. Due to the constraints of max-pooling, the model must base the bag label prediction solely on the most positive instance in the bag. If the classifier were to use $\bm{Z}_{B}$ to determine the instance label, because the bag-inherited information, such as staining conditions, is not entirely consistent across positive slides and differs on negative slides, the model would completely fail to fit the training set. Therefore, models that follow the MIL assumption must learn features such as the morphological structure of tumour cells, $\bm{Z}_{c}$, to fit the training set. A significant advantage is that models respecting MIL assumptions will achieve better patch-level classification results and enhanced out-of-distribution generalization.
Why Max-Pooling-Based MIL Methods Sometimes Underperform?
We believe there are two main reasons why the current max-pooling-based MIL methods do not perform ideally in the WSI classification task. The first reason is that mainstream methods like mi-Net (Wang et al., 2018) tend to directly remember the instances in the bag to fit the training set when the positive concept is difficult to learn. As shown in Figure 3, mi-Net perfectly fits the training set at the slide-level, while it produces almost random predictions in the validation set and fails almost completely for the patch-level predictions. This suggests that the model overfits by memorizing instances within the bag. As mentioned in Section 4.1, the reason for the difficulty in learning the positive concept is usually that the tumour region in the slide is small and tumours cells in some slides resemble normal cells. Empirically, we found that this problem disappears when using features extracted by advanced foundation models such as CTransPath (Wang et al., 2022).
The second reason, as discussed in Section 4.1, is that the malignancy levels of tumour cells across different slides are different. Since existing max-pooling-based methods select the most positive instance within one bag in a batch, existing max-pooling-based methods may perform well on salient instances but poorly on hard instances. However, this problem is also a challenge for attention-based methods.
Figure 3: Training and validation metrics over epochs for the mi-Net model on the Camelyon16 dataset.
Figure 4: Overview of our FocusMIL with 2 slides in a mini-batch for illustration. Features are extracted from patches using a pre-trained feature extractor. The instance features are encoded into latent representations via a variational encoder. The classifier $P(Y|\bm{z})$ uses the inferred latent representations $\bm{z}$ to obtain patch-level prediction scores. Max-pooling is applied to the instance predictions of each MIL bag to derive the slide-level prediction.
5 FocusMIL
To address the two aforementioned challenges, we propose FocusMIL. Figure 4 illustrates the architecture of our model. Firstly, raw input patches $\bm{p}$ are processed with pre-trained feature extractors into $\bm{x}$. Then, the variational encoder encodes $\bm{x}$ into a multivariate Gaussian distribution with a prior of $\mathcal{N}(0,I)$. Using the reparameterization trick (Kingma & Welling, 2013), latent representations $\bm{z}$ are sampled from this distribution and utilized to predict patch-level labels. The patch-level predictions are aggregated with the max-pooling operation to comply with the standard MIL assumption. Another key difference between our method and mi-Net is that we optimize multiple bags simultaneously within a mini-batch. We will explain our method in the following sections.
Addressing Overfitting with Variational Encoder
Since max-pooling leads to one instance per bag being trained in one iteration, as discussed in Section 4.3, when the positive concept is difficult to learn, the model fits the training set by remembering the instance within the bag. To prevent the classifier from directly memorizing instances, we employ a variational encoder that encodes each instance into a distribution and resample from the distribution to obtain a latent representation. Inspired by Variational Autoencoders (VAEs) (Kingma & Welling, 2013), we use a KL loss to regularize the latent space, ensuring it remains smooth and meaningful. We set the prior distribution as $p(\bm{z})=\mathcal{N}(0,\bm{I})$.
Our final loss function is:
$\mathcal{L}=-\log p\left(Y\mid\bm{z}^{*}\right)+\beta\cdot\operatorname{KL}\left[q\left(\bm{z}\mid\bm{x}\right)\|p\left(\bm{z}\right)\right]$ (6)
where $\beta$ is a hyperparameter for balancing the classifier and KL loss, $\bm{z}^{*}$ corresponds to the most positive latent representation of a bag obtained with max-pooling with respect to $p(Y|\bm{z})$.
The key difference between our method and CausalMIL (Zhang et al., 2022b) is that our approach regularizes the latent space of all instances, whereas CausalMIL only regularizes the latent space of the most positive instance per bag. The regularization approach of CausalMIL results in a large number of instances’ latent spaces remaining unconstrained, making it hard to sample meaningful representations for those instances. Our subsequent experiments also demonstrate that CausalMIL cannot effectively mitigate the simple memorization of bags on more complex datasets. Additionally, since we believe that adhering to the standard MIL assumption inherently enforces the disentangling of $\bm{Z}_{c}$ and $\bm{Z}_{B}$, our method does not require instance reconstruction.
Mini-Batch Gradient Descent for WSI Classification
For WSI classification, the morphology of tumour cells may be different across slides due to the varying malignancy levels of cancer in patients (Zhang et al., 2023; Qu et al., 2024; Jögi et al., 2012). When multiple ($n$) bags are in the mini-batch, the encoder and classifier will forward propagate multiple bags simultaneously. For each bag, max-pooling selects the most positive instance. Since the most positive patches in different tumour slides within a batch exhibit varying malignancy levels, the optimization process must consider the features of these $n$ patches, thereby learning a better classification boundary. Consider the following scenario: if the model established a classification boundary based only on salient instances, it would not be able to accurately classify slides of lower malignancy, resulting in a high loss.
Batch size $>1$ also contributes to the disentanglement of $\bm{Z}_{c}$ and $\bm{Z}_{B}$. Staining conditions vary across slides. When faced with tumour cells of different colours (e.g., tumour cell patch stained purple and tumour cell patch stained pink), the model can only rely on $\bm{Z}_{c}$ for prediction. If it depends on the staining information $\bm{Z}_{B}$ inherited from the slide, it will not be able to achieve accurate predictions (Zhang et al., 2022b).
6 Experiments
Table 1: Performance Comparison of Multiple MIL Methods using ResNet-18 and CTransPath Features. The subscripts are the corresponding $95\%$ confidence intervals. The best performance is in bold, while the second-best is underlined.
Method Performance Slide-level Patch-level
AUC ACC AUCPR F1-score
ResNet-18 ABMIL 0.8052 (0.7492, 0.8612) 0.8152 (0.7743, 0.8560) 0.1449 (0.0449, 0.2449) 0.1570 (0.0658, 0.2482)
DSMIL 0.7733 (0.7167, 0.8300) 0.7984 (0.7653, 0.8316) 0.2693 (0.0409, 0.4976) 0.2468 (0.1147, 0.3790)
TransMIL 0.8352 (0.7807, 0.8898) 0.8127 (0.7865, 0.8390)
DTFD-MIL 0.8619 (0.8530, 0.8709) 0.8032 (0.7497, 0.8567) 0.2079 (0.1592, 0.2566) 0.1640 (0.1388, 0.1891)
IBMIL 0.8442 (0.8329, 0.8555) 0.7906 (0.7465, 0.8347) 0.2587 (0.2138, 0.3037) 0.1948 (0.1499, 0.2398)
mi-Net 0.8014 (0.7674, 0.8355) 0.8158 (0.7900, 0.8416) 0.4108 (0.3466, 0.4751) 0.3844 (0.3437, 0.4251)
CausalMIL 0.8092 (0.7730, 0.8454) 0.8281 (0.8128, 0.8435) 0.4845 (0.4653, 0.5036) 0.3650 (0.3469, 0.3823)
FocusMIL 0.8706 (0.8489, 0.8923) 0.8359 (0.8206, 0.8513) 0.4382 (0.3907, 0.4857) 0.3996 (0.3787, 0.4205)
CTransPath ABMIL 0.9659 (0.9624, 0.9694) 0.9375 (0.9306, 0.9444) 0.3884 (0.3335, 0.4433) 0.3402 (0.2497, 0.4306)
DSMIL 0.9299 (0.9048, 0.9551) 0.9094 (0.8931, 0.9256) 0.6639 (0.6587, 0.6690) 0.6086 (0.5986, 0.6186)
TransMIL 0.9603 (0.9357, 0.9849) 0.9281 (0.8935, 0.9627)
DTFD-MIL 0.9739 (0.9722, 0.9756) 0.9515 (0.9472, 0.9559) 0.4535 (0.4315, 0.4756) 0.3350 (0.3072, 0.3628)
IBMIL 0.9716 (0.9691, 0.9740) 0.9562 (0.9476, 0.9649) 0.4285 (0.3977, 0.4594) 0.2927 (0.2607, 0.3247)
mi-Net 0.9757 (0.9723, 0.9791) 0.9593 (0.9550, 0.9637) 0.6156 (0.6007, 0.6305) 0.5793 (0.5542, 0.6045)
CausalMIL 0.9700 (0.9661, 0.9740) 0.9531 (0.9412, 0.9650) 0.6640 (0.6509, 0.6770) 0.6107 (0.5762, 0.6452)
FocusMIL 0.9731 (0.9725, 0.9738) 0.9609 (0.9609, 0.9609) 0.6902 (0.6821, 0.6983) 0.6893 (0.6772, 0.7014)
Table 2: Localization FROC Performance Comparison. FocusMIL (1) and FocusMIL (3) refer to the FocusMIL models trained with a batch size of 1 and 3, respectively.
Method Localization FROC
ABMIL 0.2434
DTFD-MIL 0.3013
FocusMIL (1) 0.3462
FocusMIL (3) 0.3646
6.1 Datasets, Metrics, and Baselines
Camelyon16 dataset is widely used for metastasis detection in breast cancer (Bejnordi et al., 2017a). The dataset consists of 270 training and 129 testing WSIs.
Camelyon17 dataset consists of 1,000 WSIs from five hospitals. Due to the absence of labels in the test set, we only used the 500 slides from the training set. Following the setup of AEM (Zhang et al., 2024), we designated 200 slides from the fourth and fifth hospitals as the test set to evaluate the model’s OOD generalization performance. The remaining 300 slides were split into training and validation sets with an 8:2 ratio. For a more detailed discussion of the Camelyon datasets, as well as additional OOD experiment results, please refer to the Appendix.
Semi-Synthetic Dataset To test whether existing WSI classification methods respect the standard MIL assumption, we propose the Camelyon16 Standard-MIL test dataset, inspired by the ideas in (Raff & Holt, 2023). Specifically, in the training set, we introduce poison by randomly selecting $20\%$ of the patches in the normal slides and increasing the intensity of their green channel. In the test set, we randomly select $20\%$ of the patches in tumour slides to introduce poison in the same way. A MIL model cannot legally learn to use the poison signal because it occurs only in normal slides (Raff & Holt, 2023). If a model has a training AUC $>0.5$, but a test AUC $<0.5$, it relies on poison for prediction and does not respect the MIL assumption.
Evaluation Metrics Due to the high level of class imbalance at the patch level, we report Area Under the Precision-Recall Curve (AUCPR) and F1 score for evaluation. For slide-level classification on Camelyon16 dataset, we report AUC and Accuracy as the tumour and normal classes are balanced. We report AUC and F1-score on Camelyon17 dataset, as the classes are relatively imbalanced. On Camelyon16, we also evaluate tumour region localization performance by reporting free response operating characteristic curves (FROC) (Bejnordi et al., 2017a).
Baselines We compare our method with several recently published baselines, including ABMIL (Ilse et al., 2018), DSMIL (Li et al., 2021), TransMIL (Shao et al., 2021), DFTD-MIL (Zhang et al., 2022a), IBMIL (Lin et al., 2023), mi-Net (Wang et al., 2018), and CausalMIL (Zhang et al., 2022b). The first five methods use attention mechanisms, while mi-Net and CausalMIL are based on max-pooling. To the best of our knowledge, CausalMIL has not yet been applied for WSI classification. For the Camelyon16 dataset, we adopt ResNet18 (He et al., 2016) (pretrained on ImageNet (Deng et al., 2009)) and CTransPath (Wang et al., 2022) as feature extractors. For the Camelyon17 dataset, we directly utilize the features released by AEM (Zhang et al., 2024), which include features extracted by ResNet-18 and PathGen-CLIP (Sun et al., 2024). For the semi-synthetic dataset, ResNet-18 is used to extract features.
Table 3: Performance Comparison of MIL Methods on Camelyon17 Dataset using ResNet-18 and PathGen-CLIP Features.
Method Performance Slide-level
AUC F1-score
ResNet-18 ABMIL 0.7877 (0.7505, 0.8249) 0.7004 (0.6702, 0.7306)
DSMIL 0.7069 (0.5669, 0.8469) 0.6507 (0.5826, 0.7187)
TransMIL 0.6899 (0.6187, 0.7610) 0.6200 (0.5459, 0.6941)
DTFD-MIL 0.8192 (0.7964, 0.8419) 0.7242 (0.6960, 0.7524)
IBMIL 0.8221 (0.8059, 0.8383) 0.7233 (0.7022, 0.7444)
mi-Net 0.5066 (0.5022, 0.5110) 0.2901 (0.2419, 0.3383)
CausalMIL 0.5104 (0.4615, 0.5592) 0.4905 (0.3881, 0.5929)
FocusMIL 0.8719 (0.8619, 0.8818) 0.8335 (0.8273, 0.8398)
PathGen-CLIP ABMIL 0.9264 (0.9156, 0.9372) 0.8909 (0.8880, 0.8938)
DSMIL 0.9212 (0.9035, 0.9390) 0.8904 (0.8851, 0.8957)
TransMIL 0.9387 (0.9274, 0.9500) 0.8932 (0.8871, 0.8993)
DTFD-MIL 0.9447 (0.9341, 0.9553) 0.8870 (0.8788, 0.8952)
IBMIL 0.9410 (0.9308, 0.9512) 0.8854 (0.8764, 0.8944)
mi-Net 0.9442 (0.9307, 0.9577) 0.8882 (0.8819, 0.8946)
CausalMIL 0.9537 (0.9445, 0.9628) 0.8885 (0.8799, 0.8970)
FocusMIL 0.9529 (0.9456, 0.9602) 0.8974 (0.8875, 0.9073)
6.2 Results on Real-World Dataset
Results on Camelyon16 As shown in Table 1, for patch-level classification, max-pooling-based methods such as mi-Net, CausalMIL, and FocusMIL achieve significantly better results than the attention mechanism-based MIL methods. For slide-level classification, mi-Net and CausalMIL achieve performance comparable to ABMIL. Our proposed FocusMIL achieves better results than the more sophisticated attention-based methods in terms of both AUC and ACC. For tumour region localization (Table 2), FocusMIL outperforms DTFD-MIL substantially. It is worth mentioning that after using the mini-batch gradient optimization strategy, FocusMIL achieves better small-area tumour (hard instance) prediction compared to training with SGD. We provide extensive visualization results in the Appendix.
When using features extracted by CTransPath, all models show a noticeable performance boost. In this case, the max-pooling-based methods still achieve significantly better results at the patch-level. For slide-level, three max-pooling-based methods achieve slightly leading results. Notably, DSMIL achieves better results compared to other attention-based methods. This improvement is attributed to DSMIL’s first stream using a max-pooling-based instance classifier.
Results on Camelyon17 Since the training and testing data come from different hospitals, the distribution of the bag context information $\bm{Z}_{B}$ may vary. In this OOD scenario, as shown in Table 3, FocusMIL achieves significantly better results. However, mi-Net and CausalMIL overfit the dataset by memorizing instance features within the bag, leading to training failure. After using features extracted by the advanced foundation model, the three max-pooling-based methods still exhibit slightly better performance.
6.3 Results on Standard-MIL test
As shown in Table 4, all methods fit well in the training set. All of the tested attention-based methods have a test slide AUC less than 0.5, failing to respect the standard MIL assumption. The three max-pooling-based methods are little affected by the poison. Benefiting from the max-pooling-based instance classifier, DSMIL performs well in patch-level classification, further demonstrating that max-pooling is crucial for learning positive concepts.
For the ablation study on FocusMIL, and the effect of batchsize and KL coefficients on the model performance, and computational cost analysis, please refer to Appendix.
Table 4: Performance on Camelyon16 Standard-MIL test dataset.
Model Training Testing
Slide AUC Slide AUC Patch F1
ABMIL 1.000 0.007 (0.000, 0.021) 0.142 (0.117, 0.167)
DSMIL 1.000 0.377 (0.000, 0.776) 0.312 (0.165, 0.459)
TransMIL 1.000 0.001 (0.001, 0.002) -
DTFD-MIL 1.000 0.010 (0.000, 0.034) 0.122 (0.080, 0.165)
IBMIL 1.000 0.000 (0.000, 0.001) 0.104 (0.098, 0.110)
mi-Net 1.000 0.805 (0.772, 0.838) 0.384 (0.317, 0.453)
CausalMIL 1.000 0.808 (0.785, 0.830) 0.366 (0.325, 0.407)
FocusMIL 1.000 0.868 (0.853, 0.883) 0.384 (0.366, 0.401)
7 Conclusion
In this paper, we analyze the fact that attention-based MIL methods are inherently prone to spurious correlation in WSI classification. Furthermore, we empirically investigate the limitations of existing max-pooling-based MIL methods and introduce FocusMIL, a simple yet effective max-pooling approach. Our comprehensive experiments confirm that max-pooling-based approaches can achieve superior results to the attention-based MIL approaches and exhibit significantly more robust performance in out-of-distribution scenarios. We therefore hope our work will motivate a shift from attention-based methods and a renewed focus on max-pooling-based methods to achieve more robust, interpretable, and trustworthy clinical applications.
Impact Statement
Our paper demonstrates that attention-based MIL methods are often not reliable in computer-aided diagnosis based on histopathology images. Although advanced foundation models can significantly boost their performance, attention-based aggregation approaches remain untrustworthy for high-stakes cancer diagnosis.
In real-world scenarios, the training data and deployment data often originate from different hospitals, resembling our experimental setup on the Camelyon17 dataset. In such conditions, max-pooling-based methods exhibit a significant advantage by focusing on the morphological features of cancer cells while being much less susceptible to spurious correlations. We believe that our position—refocusing on max-pooling-based MIL methods that adhere to the standard multi-instance assumption—will drive computer-aided WSI diagnostic models towards more reliable outcomes.
Furthermore, our results stress the importance of rigorous evaluation and validation across diverse scenarios before deploying WSI diagnostic models, as an unreliable model could result in catastrophic misdiagnoses in clinical settings.
Appendix A Additional Quantitative Experiments
A.1 Ablation Study
As shown in Table 5, the baseline model mi-Net is very unstable. In our experiments, we found that mi-Net may achieve high training slide-level AUC but extremely low patch-level AUCPR during training and fails completely during testing. This severe overfitting issue disappeared when the dropout was set above 0.4. However, this “training failure” reappeared when training mi-Net using mini-batch gradient descent, since establishing correct classification boundaries for salient and hard instances within a batch of multiple slides is more difficult. FocusMIL does not incorporate dropout; instead, learning the distribution helps to avoid “training failure” and achieves a slide AUC that is 1.3% higher than mi-Net with a dropout of 0.4. Further significant improvements of 7.0% in slide AUC are obtained by training multiple bags together in a mini-batch. Since our method does not rely on bag-inherited information $\bm{Z}_{B}$ for classification, a higher slide AUC actually indicates better classification boundaries for hard instances, which is also confirmed by visualization results.
Table 5: Ablation study on Camelyon16 Dataset. Dropout is set at 0.4 for the second model.
Model Mini-batch Slide AUC Patch F1
mi-Net 0.757 $\pm$ 0.122 0.298 $\pm$ 0.167
mi-Net+dropout 0.801 $\pm$ 0.027 0.384 $\pm$ 0.033
mi-Net+dropout 0.635 $\pm$ 0.133 0.248 $\pm$ 0.118
FocusMIL(SGD) 0.814 $\pm$ 0.017 0.386 $\pm$ 0.022
FocusMIL 0.871 $\pm$ 0.011 0.400 $\pm$ 0.011
A.2 Ablation Study on Hyperparameter
Ablation Study on KL Divergence Coefficient
As shown in Figure 5, we still observe training failure when $\beta=0$. However, when $\beta$ is set to 0.001, 0.01, or 0.1, the model achieves the desired performance. This further emphasizes the importance of regularizing the latent space. When $\beta=1$, the performance starts to decline, indicating that excessive regularization negatively impacts classification performance.
Figure 5: Results of ablation experiments on KL divergence coefficient and batch size. We report the mean of five experiments
Ablation Study on Batch Size
Batch size greater than 1 can significantly improve the model’s slide-level classification performance. A batch size that is too large can reduce the performance of the model, and we recommend choosing a batch size between 2 and 7.
A.3 Out-of-Distribution Generalization
To further verify that the max-pooling based MIL methods can effectively learn causal factors $\bm{Z}_{C}$, we conduct out-of-distribution (OOD) generalization experiments. Our experimental setup is consistent with CausalMIL (Zhang et al., 2022b), so for the results of the comparison methods, we directly refer to it’s paper.
As shown in Table 6, the three max-pooling-based MIL methods achieve desirable results in both the training and test sets, indicating that they successfully learn the causal factor.
Table 6: Performance on distributionally biased ColoredMNIST dataset.
Method Train Test
ERM $0.849\pm 0.002$ $0.105\pm 0.007$
ERM1 $0.848\pm 0.002$ $0.109\pm 0.005$
ERM2 $0.850\pm 0.002$ $0.101\pm 0.002$
ROBUST MIN MAX $0.843\pm 0.004$ $0.152\pm 0.025$
IRM $0.593\pm 0.044$ $0.628\pm 0.096$
IRM GAME $0.634\pm 0.011$ $0.599\pm 0.027$
iCaRL $0.706\pm 0.008$ $0.688\pm 0.007$
MIVAE $0.810\pm 0.005$ $0.156\pm 0.003$
mi-Net $0.930\pm 0.005$ $0.914\pm 0.008$
CausalMIL $0.919\pm 0.002$ $0.892\pm 0.002$
FocusMIL $0.956\pm 0.010$ $0.935\pm 0.007$
Table 7: Comparison of model size and training time per epoch of different MIL methods. The input dimension is set the same for each model to ensure fairness of comparison.
Model Para. Time C16 AUC C17 AUC
ABMIL 657k 3.35s 0.8052 0.7877
DSMIL 856k 4.92s 0.7733 0.7069
TransMIL 2.66M 9.56s 0.8352 0.6899
DTFD-MIL 987k 6.02s 0.8619 0.8192
CausalMIL 302k 3.28s 0.8092 0.5104
FocusMIL 167k 3.23s 0.8706 0.8719
Figure 6: Visualization results of FocusMIL and DTFD-MIL on Camelyon16 dataset. Ours(3) and Ours(1) refer to the FocusMIL models trained with a batch size of 3 and 1, respectively. The fourth slide is negative. For FocusMIL, classifier confidence scores are utilized for patch intensities. For DTFD-MIL, patch probability derivation is used with a threshold of 0.5001.
Appendix B Computational Cost Analysis
As shown in Table 7, FocusMIL has significantly fewer parameters compared to attention-based methods. Due to the use of mini-batch gradient descent with a batch size of 3, both CausalMIL and FocusMIL achieve superior training speeds. However, CausalMIL requires additional computations for reconstructing loss and the KL divergence loss of the conditional prior distribution, which results in slightly longer training times compared to FocusMIL.
Appendix C Visualization
For fairness of comparison, the models for visualization are not deliberately selected, and both FocusMIL with a train- ing batch size of 3 and DTFD-MIL models achieved a test slide AUC of approximately 0.87. As shown in Figure 6, our model accurately identifies both large and small tumour regions without any false positive predictions. FocusMIL with a training batch size of 1 misses some small tumour regions. For DTFD-MIL, it misses the tumour region in the first slide and some tumour patches in the second slide, and there are many false positive predictions.
Notably, in the fourth slide of Figure 6 and the normal slide in Figure 7, DTFD-MIL assigns high scores to many patches unrelated to the tumour. After zooming in on these patches, we observe that the tissues in these regions are significantly blurrier, which is associated with the slide preparation process. This indicates that DTFD-MIL relies on bag-inherited information $\bm{Z}_{B}$ for predictions. FocusMIL does not have any false-positive predictions for these normal slides.
Figure 8 demonstrates the visualization results of FocusMIL and DTFD-MIL for five tumour slides. FocusMIL with a training batch size of 3 perfectly predicts all tumour regions. When the training batch size was changed to 1, FocusMIL misses some small tumour regions. This result confirms that training with mini-batch gradient descent does help to establish better classification boundaries for salient and hard instances. For DTFD-MIL, its predictions for both large and small tumour areas are suboptimal due to the fact that it can overfit on salient instances.
Figure 7: Visualization results of FocusMIL and DTFD-MIL on 3 normal slides. In the zoomed-in areas, some regions are relatively clear, while the rest are very blurred. DTFD-MIL may assign a high positive probability to the blurred regions.
Figure 8: Additional visualization results of FocusMIL and DTFD-MIL on Camelyon16 dataset. Ours(3) and Ours(1) refer to the FocusMIL models trained with a batch size of 3 and 1, respectively. The tumour regions are delineated by the blue lines.
Appendix D Experimental setup
D.1 Dataset and Pre-processing
Camelyon16 dataset is widely used for metastasis detection in breast cancer (Bejnordi et al., 2017a). The dataset consists of 270 training and 129 testing WSIs, which yield roughly 2.7 million patches at 10× magnification (the second level in multi-resolution pyramid). Pixel-level labels are available for tumour slides. Every WSI is cropped into $256\times 256$ patches without overlap, and background patches are discarded. A patch is labeled positive if it contains $25\%$ or more cancer areas. The numbers of tumorous versus normal patches are imbalanced as the positive patch ratios of positive slide in the training and testing sets of Camelyon16 are approximately 8.8% and 12.7%, respectively. The complete code for pre-processing will be made publicly available.
Camelyon17 dataset consists of 1,000 WSIs from five hospitals, categorized into different slide labels such as Normal, Isolated Tumour Cells, Micro-metastases, and Macrometastases (Bejnordi et al., 2017b). Since this study focuses on a binary classification task, we relabeled the slides into two categories: Normal and Tumour. Due to the absence of labels in the test set, we only used the 500 slides from the training set. Following the setup of AEM (Zhang et al., 2024), to evaluate the model’s OOD generalization performance, we designated 200 slides from the fourth and fifth hospitals as the test set. The remaining 300 slides were split into training and validation sets with an 8:2 ratio.
D.2 Implementation details
The encoder of FocusMIL consists of a neural network with one hidden layer, with ReLU as the activation function. We use AdamW optimizer with an initial learning rate of $1\times 10^{-4}$ to update the model weights during the training. The dimension of the latent factor is set to 35. The mini-batch size for training FocusMIL model is 3. All experiments are conducted 5 times, and the mean and 0.95 confidence intervals (CI) are reported.
For max-pooling-based methods, we directly use the classifier output as the patch-level prediction. For ABMIL, following the DTFD paper (Zhang et al., 2022a), we use the normalized attention scores for patch-level prediction. For DTFD-MIL, we use the instance probability derivation proposed in their paper for patch-level prediction.
For the attention-based methods, we build other models based on their officially released codes and conduct grid searches for key hyperparameters. For ABMIL, we use gated attention with dropout rate of 0.25 to achieve optimal performance. For CausalMIL (Zhang et al., 2022b), both its encoder and decoder are set up as neural networks with hidden layer neurons 128 and two other fully-connected layers are used to carve factorized prior distribution conditioned on the bag information. All the experiments are conducted with a Nvidia RTX4090 GPU. For more details, please refer to our code in the supplementary.
Appendix E Future Work
In this paper, we have demonstrated that max-pooling-based methods, which adhere to the standard multi-instance assumption, can yield more robust results for WSI classification task. Here are several directions for future exploration:
  1. 1.
    Beyond WSI classification. There may be other tasks that inherently conform to the standard multi-instance assumption but still widely adopt MIL methods that violate this assumption. Re-visiting max-pooling-based approaches in those domains could contribute to more robust decision-making.
  2. 2.
    Alleviating overfitting. Exploring alternative strategies to mitigate overfitting in max-pooling-based methods is a promising direction. Additionally, investigating more complex priors beyond multivariate Gaussian distributions is also worth exploring as a potential avenue for improvement.
  3. 3.
    Improving detection of low-malignancy tumour regions. Identifying small, low-malignancy cancerous areas remains a crucial and challenging problem. While this paper introduces mini-batch optimization, other approaches are also worth exploring.